OpenStudy (anonymous):

i need urgent help...:prove that the projection of vector A onto vector B is equal to vector A * the unit vector of B

OpenStudy (anonymous):

any further info on question?

OpenStudy (ybarrap):

Projection of A onto B = A*B (where * is vector dot product) Let |*| is the magnitude of the vector * A*B = |A||B|* cos (x) , where x is the angle between A and B A*B/|B| = |A|*cos(x) Now |A|cos(x) is the projection of A onto B And B/|B| is the unit vector and we're done.